Let the country of Coalistan have 7 coal producers, cleverly named A-G. Let A have 30%...

  1. Let the country of Coalistan have 7 coal producers, cleverly named A-G. Let A have 30% of the market, B have 20% of the market, and C through G each have 10% of the market. What is the herfindahl index for this market?

    .09

    .10

    .18

    .20

    None of the above
Answer #1
Producer Market share Square of market share
A 0.30 0.09
B 0.20 0.04
C 0.10 0.01
D 0.10 0.01
E 0.10 0.01
F 0.10 0.01
G 0.10 0.01
- Sum of squares of market shares 0.18

Herfindahl Hirschman Index is obtained as the sum of squares of market shares of all the firms in the market.

HHI = Sum of squares of market shares of all the firms = 0.3^2 + 0.2^2 + 0.1^2 + 0.1^2 + 0.1^2 + 0.1^2 + 0.1^2 = 0.09 + 0.04 + 0.01 + 0.01 + 0.01 + 0.01 + 0.01 = 0.18

Ans: 0.18